OpenStudy (anonymous):

Let R be the region between the graph of f and the x-axis on the interval . Find the volume V of the solid of revolution generated by revolving R about the x-axis. Give both exact (symbolic) and approximate (numeric) answers f(x)=8x-x^2

OpenStudy (anonymous):

so to find this i squared the function, multiplied it by pi, and found the indefinite integral of this from 0 to 8

OpenStudy (anonymous):

the from 0 to 8 did not end up here when i copied and pasted it sorry

OpenStudy (anonymous):

so i got pi(8x^2-x^4) then i got 8pix^2-pix^4

OpenStudy (anonymous):

the indefinite integral i got from that is (8/3)pix^3 - (pi/5)x^5

OpenStudy (anonymous):

i then plugged in 8 and plugged in 0 and subtracted 0 from 8 so i got [(8pi/3)8^3 - (pi/5)8^5] - [(8pi/3)0^3 - (pi/5)0^5] which gets simplified to (8pi/3)8^3 - (pi/5)8^5 which gets simplified to (4096pi/3) -(32768pi/5)

OpenStudy (anonymous):

but its wrong

OpenStudy (anonymous):

(8x-x^2)^2 is not 8x^2-x^4

OpenStudy (anonymous):

its not?

OpenStudy (anonymous):

(a+b)^2 = a^2 +2ab +b^2

OpenStudy (anonymous):

oh i have to foil?

OpenStudy (anonymous):

so it is 64x^4-16x^6+x^8?

OpenStudy (sirm3d):

\[\large (8x-x^2)^2=64x^2-16x^3+x^4\]

OpenStudy (anonymous):

you dont add the exponents? when multiplying

OpenStudy (anonymous):

so the antiderivative is (64*pi/3)x^3 - (16*pi/4)x^4 +(pi/5)x^5 correct?

OpenStudy (anonymous):

yes

OpenStudy (sirm3d):

100% correct

OpenStudy (anonymous):

so when i put that equation in just plug in 8 for all the x why is it wrong?

OpenStudy (anonymous):

(64*pi/3)8**3-(16*pi/4)8**4+(pi/5)8**5 i put that in to be exact

OpenStudy (anonymous):

the ** represents exponents

OpenStudy (anonymous):

since when you plug in 0 to x it equals 0 i did not subtract it, just plugged in 8 for x..correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so do you know why it marked me wrong?

OpenStudy (anonymous):

i have 1 more chance

OpenStudy (anonymous):

no.

OpenStudy (anonymous):

probably calculation error

OpenStudy (anonymous):

i didnt calculate it i plugged it in like this (64*pi/3)8**3-(16*pi/4)8**4+(pi/5)8**5

OpenStudy (anonymous):

but its ok i used wolfram so i dont know how to do it but i got credit for it on my hw thx guys!